OpenStudy (anonymous):

The volume of a gas is 250 mL at 3.5 atm pressure. What will the volume be when the pressure is reduced to 0.75 atm. assuming the temperature remains constant?

OpenStudy (perl):

we have three gas laws 1. PV = k (temperature held constant) 2. V/T = k (pressure held constant) 3. P/T= k (volume held constant) And the combined gas law is PV = kT ANd the mole gas law is PV = nRT we assume ideal gases http://chemistry.bd.psu.edu/jircitano/gases.html

OpenStudy (anonymous):

I got 1166.67 mL

OpenStudy (perl):

P1*V1 = k P2*V2 = k since it is the same temperature P1 *V1 = P2*V2 (3.5 atm) ( 250 ml) = (.75 ml) (x) So x = 1166.67